Z follows a standard normal distribution, What is the probability of getting P ( Z > 0.15 ) ? Enter your answer to 4 decimal places
Answer #1

Answer

We have Pz 0.15)

Using the identity P(z>a) =1-P(z<a)

so, we can write it as P(z>0.15) =1-P(z<0.15)

using z distribution table, we get Pz < 0.15 = 0.5596 (looking 0 in the left most column and 0.15 in top most row)

setting the values, we get

P(z > 0.15) = 1-P(z < 0.15) = 1-0.5596 = 0.4404

So, required probability is 0.4404 (rounded to 4 decimals)